OpenStudy (anonymous):

Medal! The chart shows the membership in the High School Band. What is the probability that a boy and a girl chosen randomly will both be seniors? HS Band Freshman Sophmores juniors seniors Boys 10 7 10 9 Girls 8 11 9 7 A. 1/4 B. 1/5 C. 1/9 D. 1/20

OpenStudy (anonymous):

i got B as my answer

Directrix (directrix):

I wonder if the order of selection matters. That is, BG and then GB. Somehow, I don't think so. @Yttrium What do you think?

OpenStudy (yttrium):

What I know is that it is all about total probability theorem.

Directrix (directrix):

P(SB) = 9/36; P(SG) = 7/35 P(SB and SG) = (9/36)*(7/35) = ?

OpenStudy (agent0smith):

Order doesn't appear to matter - you know you're choosing from the boys, then you know you're choosing from the girls HS Band Freshman Sophmores juniors seniors Boys 10 7 10 9 total boys = 36 Girls 8 11 9 7 total girls = 35 senior boy = 9/36 = 1/4 senior girl = 7/35 = 1/5 multiply those two.

OpenStudy (anonymous):

multiply which two?

OpenStudy (anonymous):

oh i see now.

OpenStudy (agent0smith):

the only two that are there... the same two directrix has

OpenStudy (anonymous):

1/20

Directrix (directrix):

@kropot72 Please tell us what you think.

OpenStudy (anonymous):

is 1/20 the right answer guys?

OpenStudy (agent0smith):

@Directrix it's essentially conditional. "boy and a girl chosen randomly" = Given you pick a boy, given you pick a girl. You're choosing a boy from the group of boys, you're not choosing from the entire group. Same for the girl.

OpenStudy (agent0smith):

It doesn't mean the same as "what is the probability you pick a boy and a girl and both are seniors"

OpenStudy (agent0smith):

But probability problems often have these annoying wording issues...

OpenStudy (kropot72):

The events 'choose a boy' and 'choose a girl' are independent. Therefore the required probability is P(pick a senior girl) * P(pick a senior boy) = (9/36) * (7/35) = 1/20